OpenStudy (anonymous):

how do you solve this: a ball falls to the ground at a rate of sqrt(2h/g) if dropped from 10m how long does it take to come to rest. g=9.8 ans:sqrt(20/g)(1+sqrt(p))/(1-sqrt(p))

OpenStudy (nottim):

Do you use the 5 Kinematic equations here?

OpenStudy (anonymous):

not a physics problem. calc2.

OpenStudy (anonymous):

@experimentX

OpenStudy (nottim):

oh. :(

OpenStudy (anonymous):

ya, that's definitely my face!

OpenStudy (anonymous):

Integrate the equation in h from ten to zero

OpenStudy (anonymous):

where does the (1+sqrt(p))/(1-sqrt(p)) come in though?

OpenStudy (experimentx):

perhaps sqrt(2hg)

OpenStudy (anonymous):

the book says sqrt(2h/g)

OpenStudy (experimentx):

ok ok ... what is p?

OpenStudy (anonymous):

it doesn't say!!!

OpenStudy (anonymous):

Here's the equation: \[\int\limits_{10}^{0} \sqrt{2h \div g}\]

OpenStudy (experimentx):

the velocity is given as a function of distance. find the average velocity.

OpenStudy (anonymous):

Let me put in g.\[\int\limits_{10}^{0}\sqrt{h \div 4.9}\]

OpenStudy (anonymous):

That's simplifying the 2/9.8

OpenStudy (experimentx):

divide the whole integration by 10. that will give you average velocity. \[ \text{avg vel = }\frac{1}{10} \int_0^{10} \sqrt{\frac{h}{4.9}}dh\]

OpenStudy (anonymous):

That doesn't really do anything.

OpenStudy (anonymous):

i get that numb3r1, i need to find out what p is doing in this equation.

OpenStudy (anonymous):

Let's walk through the integration instead.

OpenStudy (anonymous):

you both do it your own ways.

OpenStudy (experimentx):

looks like the answer is what numb3r1 got

OpenStudy (anonymous):

i get how to integrate, this i already did that.

OpenStudy (anonymous):

.451754sqrt(h)

OpenStudy (anonymous):

\[\int\limits_{10}^{0} \sqrt{\frac{ h }{ 4.9 }}dh\] That should not be in the format sqrt (h). The integral of sqrt(h) dh is in sqrt (h^3)

OpenStudy (anonymous):

and please excuse my pissed off behavior it's 314 in the morning.

OpenStudy (anonymous):

No worries, it's past midnight here too.

OpenStudy (anonymous):

OK, I'll just do the integral.

OpenStudy (nottim):

Ha. Sleep. I use to go on 4 hr sleep a day for a yr.

OpenStudy (experimentx):

there shouldn't be 'p' ... perhaps you missed something in the question or typo on answer.

OpenStudy (anonymous):

\[\int\limits_{10}^{0}\sqrt{\frac{ h }{4.9 }} dh={\frac{ 1 }{4.9 }} \int\limits_{10}^{0}\sqrt{h}\]

OpenStudy (anonymous):

Sorry, that's wrong.

OpenStudy (anonymous):

It should be root(4.9)

OpenStudy (anonymous):

what the?? can you do that?

OpenStudy (anonymous):

AH. better.

OpenStudy (anonymous):

The root(4.9) is just a constant. Constants can be put in or taken out of integrals

OpenStudy (anonymous):

sure, i know that much.

OpenStudy (anonymous):

2x^(3/2)/3

OpenStudy (anonymous):

h is x

OpenStudy (anonymous):

\[\int\limits_{10}^{0}\sqrt{h}dh=\int\limits_{10}^{0}h^{1/2} dh\]

OpenStudy (anonymous):

i've already gone ahead of you.

OpenStudy (anonymous):

So yes, 2x^(3/2)/3 (then put back the 4.9)

OpenStudy (anonymous):

Erm... root 4.9

OpenStudy (anonymous):

Substitute in 10 for h, then subtract the same thing with 0.

OpenStudy (anonymous):

root 1/4.9

OpenStudy (anonymous):

Yeah, that.

OpenStudy (anonymous):

sure, easy enough.

OpenStudy (anonymous):

OK, what does that get you?

OpenStudy (anonymous):

9.523

OpenStudy (anonymous):

appx

OpenStudy (anonymous):

OK, let's try not simplifying out g. In that case, you extract (sqrt 2/g) from the equation.

OpenStudy (anonymous):

sure. that leaves sqrt2/g (21.082) appx

OpenStudy (anonymous):

Can you check the answer to make sure you typed it in correctly?

OpenStudy (anonymous):

and yes it's all correct.

OpenStudy (anonymous):

Solve for p and see if it's anything useful?

OpenStudy (anonymous):

huh???

OpenStudy (agent0smith):

This question... calculus problem or not, something seems off. a rate of sqrt(2h/g) hmm, rate implies a velocity in m/s... but that equation does not give you velocity, it gives you time for an object to fall. You can see here: http://www.gravitycalc.com/ \[\Large t = \sqrt { \frac{ 2h }{ g } }\] You can confirm this by looking at the units of each (h is in metres, g is m/s^2): \[\Large t = \sqrt { \frac{ m }{m s^{-2} } } = \sqrt{s^2} = seconds\] Velocity IS given by \[\Large v = \sqrt { 2g h }\] And this too "how long does it take to come to rest" ... I guess that's supposed to mean "how long does it take to hit the ground"? Or are we supposed to assume it bounces when it hits the ground, and find the time it takes to stop bouncing...?

OpenStudy (anonymous):

it bounces.

OpenStudy (anonymous):

trust me i've been doing this problem for about 2 weeks, and it still doesn't make any sense.

OpenStudy (agent0smith):

Well that seems like the fault of the question, it's a poorly worded/constructed problem... I have to guess p is a measure of the springyness of the ball.

OpenStudy (john_es):

As @agent0smith says the rate that the problem states is really the time to hit the ground. The problem is poorly constructed.

OpenStudy (anonymous):

yes, i would have to agree with you there.

OpenStudy (anonymous):

well, i paraphrased the question but there's no mention of p anywhere in the problem.

OpenStudy (agent0smith):

 If those are the bounces, we have to find a way to sum up the time for each one. http://www.sosmath.com/calculus/geoser/bounce/bounce.html